1. The demand for a product, in dollars, is = 2000 - 0.2- 0.01x2

Find the consumer surplus when the sales level is 250.

2.1. Find the area of the region that lies inside the circle r = 3sin(x)and

outside the cardioid r = 1+ sin(x).

2.2 Find the length of the cardioid r = 1+ sin(x)


1
Expert's answer
2022-05-04T14:51:26-0400

1. Since the number of products sold is "x=250," the coresponding price is


"P=2000-0.2(250)-0.01(250)^2=1325"

The consumer surplus is as follows


"\displaystyle\int_{0}^{250}(p(x)-P)dx"

"=\displaystyle\int_{0}^{250}(2000-0.2x-0.01x^2-1325)dx"

"=[675x-0.1x^2-\dfrac{0.01x^3}{3}]\begin{matrix}
 250 \\
 0
\end{matrix}"

"=675(250)-0.1(250)^2-\dfrac{0.01(250)^3}{3}"

"=\$110416.67"

2.1


The cardioid intersects with the circle


"3\sin\theta=1+\sin \theta"

"\sin \theta=\dfrac{1}{2}"

The cardioid intersects with the circle at "(\dfrac{3}{2},\dfrac{\pi}{6}),(\dfrac{3}{2},\dfrac{5\pi}{6})" and the pole.

The area of interest has been shaded above.

To find the area of a polar curve, we use


"A=\dfrac{1}{2}\displaystyle\int_{\pi/6}^{5\pi/6}((3\sin \theta)^2-(1+\sin \theta)^2)d\theta"

"=\dfrac{1}{2}\displaystyle\int_{\pi/6}^{5\pi/6}(8\sin^2 \theta-2\sin\theta-1)d\theta"

"=\dfrac{1}{2}\displaystyle\int_{\pi/6}^{5\pi/6}(3-4\cos2 \theta-2\sin\theta)d\theta"

"=\dfrac{1}{2}[3\theta-2\sin2\theta+2\cos \theta]\begin{matrix}
 5\pi/6\\
 \pi/6
\end{matrix}"

"=\dfrac{1}{2}(\dfrac{5\pi}{2}+\sqrt{3}-\sqrt{3}-\dfrac{\pi}{2}+\sqrt{3}-\sqrt{3})"

"=\pi"

"\pi" square units.


2. 2

"r(x)=1+\sin x, r'(x)=\cos x"

"(r)^2+(r')^2=(1+\sin x)^2+(\cos x)^2"

"=2+2\sin x"

"L=\displaystyle\int_{-\pi/2}^{3\pi/2}\sqrt{(r(x))^2+(r'(x))^2}dx"

"=\displaystyle\int_{-\pi/2}^{3\pi/2}\sqrt{2+2\sin x}dx"

"=2\displaystyle\int_{-\pi/2}^{3\pi/2}|\cos(\dfrac{x}{2}-\dfrac{\pi}{4})|dx"

"=4\bigg[\sin(\dfrac{x}{2}-\dfrac{\pi}{4})\bigg]\begin{matrix}
 3\pi/2 \\
 -\pi/2
\end{matrix}=4(1+1)=8(units)"

8 units